OpenStudy (anonymous):

Write the sum using summation notation, assuming the suggested pattern continues. -4 + 5 + 14 + 23 + ... + 131

OpenStudy (anonymous):

count the spaces between each of the numbers, the spaces are 9, so count from 23 and add 9 and that you answer understand?

OpenStudy (anonymous):

whats 23 plus 9

OpenStudy (anonymous):

32??

OpenStudy (anonymous):

yes :D

OpenStudy (jack1):

think ur missing the point of summation notation http://www.columbia.edu/itc/sipa/math/summation.html

OpenStudy (anonymous):

thanks @callie2240

OpenStudy (anonymous):

this are the options i have @callie2240 \[\sum_{n=D}^{15} (-4+9n)\] \[\sum_{n=D}^{15} -36n\] \[\sum_{n=D}^{\infty} (-4+9n)\] \[\sum_{n=D}^{\infty} -36n\]

OpenStudy (anonymous):

forget the last two, you are only summing a finite number of terms

OpenStudy (anonymous):

unless there are \(...\) after the 131

OpenStudy (anonymous):

i assume the first one is actually \[\large \sum_{n=0}^{15} (-4+9n)\]

OpenStudy (anonymous):

ok but why?